2015 Paper (5 marks) 5. Third order system Solve the differential equation x + 14x + 61x + 74x = 74 with initial conditio
Answer #1

egun Man we sove these G+C2 = 4 24 + 6C2-3 = 7 46 + 35C - 12 = 4. put = 4 - in f 1 me eget. 8-24 + 6 - 3 = 7. 16-46₂ + 3 C -